a hall has dimensions 10m*12m*14*.a fly stating at one corner ends up to a diaetrically opposite corner.the magnitude of its displacement is nearly

Grade:12

1 Answers

light yagami
6 years ago
since teh ans content should not be less than 100 chars .. i am blah blah …..................................................................................................................................displacement = sqrt(lbh) 
therefore displacement = sqrt(10*12*14) = sqrt(1680)= 40.98